if a quadratic trinomial cannot be factored, what kind of trinomial is it? Give example and explain.
Answer by maali40(13)

Quadractic is polynomial of degree 2.A1*x^2+A2*x+A3 some like Ax^2 +Bx + C Notice (x-3)*(x-4) not= x^2 -7*X +12 'WHEN CAN YOU GET AWAY WITH SAYING IT IS????' (-3)*(-4) = (-1)^2*12 What's (-1)^2? When? What's (x^2-4)/(x-2) are u at level u "STAY AWAY FROM 'DIVISION BY 1-1"? wHY? cAN'T GENERALLY SOLVE FOR x 'complete square heh both work the "roots" are ,when have x^2+bx +c ,(-b+(or-)(b^2-4*c)^.5)/2 where b^2=>4*c OR NO ROOTS UNLESS SOME FANTASY!!DID YOU SEE SOMEWHERE 'DONE WITH FACTORING THEN PROCEEDS MORE"FACTORING" AND "FORMULAS" What did u think about that???Are u studying algebra for the 'grade' ? then stick it